OpenStudy (anonymous):

MEDAL <3 A jug of milk holds 4 quarts. How many liters of milk does the jug hold? Round your answer to the nearest hundredth of a liter. 1 L= 1.06 qt

OpenStudy (anonymous):

@Musics_life @humphriescody please help

OpenStudy (shinalcantara):

\[4 qt (\frac{1L}{1.06qt}) = ?\]

OpenStudy (anonymous):

you take 1.06 x 4 = 4. 3 liter

OpenStudy (anonymous):

do you get it or do you need more exsplanashin

OpenStudy (shinalcantara):

just by looking at it using the units, it won't be the answer @humphriescody take note that 1.06 is in quarts so does 4..

OpenStudy (anonymous):

wait I am lost so what is your answer @humphriescody @shinalcantara

OpenStudy (anonymous):

@confluxepic @midhun.madhu1987 please help

OpenStudy (shinalcantara):

Take a look at this: if you have 1000 cm and you are to convert it into meters 1m = 100 cm \[1000cm(\frac{ 1m }{ 100cm }) = 10m\] cm will cancel out and meters will be left as a unit. ---------------------- Applying the same concept to the problem, you'll have: \[4qt (\frac{ 1L }{ 1.06qt }) = 3.77L\]

OpenStudy (anonymous):

thx!